# Risk Twin: Real-time Risk Visualization and Control for Structural Systems

### v1

@@ Appendix A Implementation details of the mechanical arm control
1. a. Read parameters, $\theta_{i}$, from mechanical arm with current position;
— b. Compute the current position $\boldsymbol{p}=(u^{0},v^{0})$ of the endpoint of the mechanical arm;
— c. Define the number of discretization $n_{\tau}$;
— d. Define $\boldsymbol{p}_{c}=(u_{c},v_{c})$.
2. Compute $\Delta{\boldsymbol{p}}=(\boldsymbol{p}_{c}-\boldsymbol{p})/n_{\tau}$;
3. If the mechanical arm touches the beam:
— a. Yes, compute $m_{i}$ using the controlled phase equation;
— b. No, compute $m_{i}$ using the uncontrolled phase equation.
4. $\boldsymbol{p}\leftarrow\boldsymbol{p}+\Delta{\boldsymbol{p}}$ and estimate $\delta{\boldsymbol{\theta}}^{\star}$ based on Eq. (25);
5. Control the arm moving to $\boldsymbol{p}$ using $\delta{\boldsymbol{\theta}}^{\star}$;
6. If $\boldsymbol{p}$ reaches $\boldsymbol{p}_{c}$:
— a. Yes, stop;
— b. No, go back to step 3.

### v2

@@ Appendix A Implementation details of the mechanical arm and wind turbine control
1:a. Read parameters, $\theta_{i}$, from the mechanical arm with the current position;
2:b. Compute the current position $\boldsymbol{p}=(u^{0},v^{0})$ of the endpoint of the mechanical arm;
3:c. Define the number of discretizations $n_{\tau}$;
4:d. Define $\boldsymbol{p}_{c}=(u_{c},v_{c})$.
5:Compute $\Delta{\boldsymbol{p}}=(\boldsymbol{p}_{c}-\boldsymbol{p})/n_{\tau}$;
6:if the mechanical arm touches the beam then
7:     a. Yes, compute $m_{i}$ using the controlled phase equation;
8:     b. No, compute $m_{i}$ using the uncontrolled phase equation.
9:end if
10:$\boldsymbol{p}\leftarrow\boldsymbol{p}+\Delta{\boldsymbol{p}}$ and estimate $\delta{\boldsymbol{\theta}}^{\star}$ based on Eq. (25);
11:Control the arm moving to $\boldsymbol{p}$ using $\delta{\boldsymbol{\theta}}^{\star}$;
12:if $\boldsymbol{p}$ reaches $\boldsymbol{p}_{c}$ then
13:     a. Yes, stop;
14:     b. No, go back to step 3.
15:end if
Algorithm 2 Reliability-informed control of the wind turbine
1:a. Read parameters from Risk Twin;
2:b. Define the control parameters $\Delta\theta$ and $\beta_{thr}$.
3:Define the proposed decisions, $D_{i}$:
4:    $\theta_{Y}(D_{i})\leftarrow\theta_{Y}+m\cdot\Delta\theta$,
5:    $\theta_{p}(D_{i})\leftarrow\theta_{p}+n\cdot\Delta\theta$, where $m,\ n\in[-1,0,1]$;
6:Estimate the prior reliability $\hat{\beta}(D_{i})$ and power $\hat{P}(D_{i})$ of the $D_{i}$:
7:    a. $\hat{\beta}(D_{i})=\min{[\beta_{blade}(D_{i}),\beta_{hub}(D_{i}),\hat{\beta}_{tower}(D_{i})]}$.
8:    b. $\hat{P}(D_{i})=Power[D_{i}]$.
9:if $\hat{\beta}(D_{i})\geq\beta_{thr}$ then
10:     a. Yes, select decision $D_{i}$ that maximizes $\hat{P}(D_{i})$, go back to step 1;
11:else
12:     b. No, select decision $D_{i}$ that maximizes $\hat{\beta}(D_{i})$, go back to step 1.
13:end if